### v1

# Direct and inverse abstract Cauchy problems with fractional powers of almost sectorial operators

Abstract.
We show the explicit solution operator of an abstract Cauchy problem involving a time-variable coefficient and a fractional power of an almost sectorial operator. Also, by using the representation of the solution operator, we solve the inverse abstract Cauchy problem for recovering the time-variable coefficient. As a complement, we also study similar problems by using almost sectorial operators that depend on a time-variable.
Key words and phrases:
Direct abstract Cauchy problems, inverse abstract Cauchy problems, almost sectorial operators, time-variable coefficient, bounded and unbounded operators, dense and non-dense domains.
2010 Mathematics Subject Classification:
65J08; 65N21, 34K08, 47A05.
1. Introduction
In functional analysis, specifically in operator theory, the Hille–Yosida theorem characterizes the generators of strongly continuous semigroups of linear operators on Banach spaces. This characterization requests operators to be closed and densely defined. Moreover, the following estimate holds:
$\|(\lambda I-A)^{n}\|\leqslant\frac{C}{(\lambda-\omega)^{n}},\quad\lambda>\omega,$
for all positive integers $n$ and any $\lambda$ in the resolvent set $\rho(A).$ Note that for contraction semigroups, the Hille–Yosida theorem can be easily characterized by those operators $A$ such that $\|(\lambda I-A)^{-1}\|\leqslant\frac{1}{\lambda},$ for any $\lambda\in\rho(A)$ with $\lambda>0.$ Nevertheless, there are different classes of important and interesting operators who do not satisfy the above estimates. Indeed, we recall the so-called almost sectorial operators, see Definition 2.1. More details about this type of operators can be found e.g. in [7, 14, 18]. It is known that operators satisfying condition (b) do not generate a $C_{0}$-semigroup since it does not satisfy the Hille-Yosida’s estimates. But, in this case, the operator $A$ will generate a type of semigroup called analytic semigroup of growth oder $\gamma$ (see Definition 2.2). This concept was first introduced by Da Prato in [5] for positive integer orders. The generalization for any positive order was given in different works e.g. [10, 11, 18, 20] and the references therein. These semigroups do not imply strong continuity at $t=0,$ and this is one of the main difference compare with the $C_{0}$-semigroups. Also, in this work, the operators do not need to have dense domains and/or ranges. This generality is not always assumed everywhere since some researchers studied almost sectorial operators defined over domains that are dense as well, see e.g. [2]. Almost sectorial operators frequently appear by the consideration of elliptic operators in regular spaces. Let us now recall briefly some classical examples in this setting. For instance, the minus Laplacian in a bounded domain $\Omega$ is sectorial (this means $\gamma=1$ in Definition 2.1) under some suitable boundary conditions in $L^{p}(\Omega)$ [4, Section 1.3]. Moreover, it is also sectorial in the spaces of bounded or continuous functions [6, 17]. While, in the space of Hölder continuous, functions the minus Laplacian is almost sectorial, see e.g. [19] or [9, Example 3.1.33]. Some other good examples of almost sectorial operators can be found in [14, Section 2]. There is an interesting example of an operator who involves Riemann-Liouville fractional derivatives in the Banach space $C_{\infty}(\mathbb{R};\mathbb{C}^{2})$ of continuous functions which vanish at infinity [15, Example 3.2], that is not sectorial but it is almost sectorial. In [15], more examples of almost sectorial operators are given by using Riemann-Liouville derivatives.
In this paper, we show explicitly the unique solution operator for the abstract Cauchy problem involving a time-dependent variable coefficient and a fractional power of an almost sectorial operator. By using the representation of the solution operator, we then consider two direct abstract Cauchy problems with the same time-variable coefficient, with the target of finding the time-dependent coefficient (inverse abstract Cauchy problems). Here, we mainly follow some ideas from [12] and [14].
We complement our studies, by considering the case of non-autonomous evolution equations with almost sectorial operators. Here, the operator depends on a time-variable. For some works in this direction, we recommend to check e.g. [3, 8]. Again, for this case, we find the unique solution operator (explicitly) for the abstract Cauchy problem with a time-dependent variable coefficient and an almost sectorial operator $A(\tau)$ for $\tau\geqslant 0$. Also, the inverse abstract Cauchy problem with this type of operators is studied.
The structure of the paper is as follows. In Section 2, we recall some definitions and results about almost sectorial operators and abstract evolution equations. We continue with Section 3, where we give some of the main results on the representation of the solution operator of abstract Cauchy problems with time-variable coefficients. While, in Section 4, we consider the inverse abstract Cauchy problem in finding the time-dependent variable coefficient. In the end (Section 5), we study similar problems as mentioned before by means of almost sectorial operators that depends on a time-variable.
2. Preliminaries
Everywhere below, we assume that $(X,\|\cdot\|)$ is a complex Banach space. Here, an operator in $X$ means a linear map $A:D(A)\subset X\to X$ whose domain $D(A)$ is a linear subspace of $X.$ Also, $\sigma(A)$ is the spectrum of $A,$ and $R(A)$ is the range. Moreover, $\mathscr{L}(X)$ denotes the space of all bounded linear operators on $X.$
Let $0<\mu<\pi$. So, we define by $S_{\mu}^{0}$ to be the open sector $\{z\in\mathbb{C}\setminus\{0\}:|\text{arg}\,z|<\mu\},$ and its closure $S_{\mu}:=\{z\in\mathbb{C}\setminus\{0\}:|\text{arg}\,z|\leqslant\mu\}\cup\{0\}.$ We consider the function arg with values in $(-\pi,\pi].$
Definition 2.1.
Let $-1<\gamma<0$ and $0\leqslant\omega<\pi.$ By $\Theta_{\omega}^{\gamma}(X)$ we denote the set of all closed linear operators $A:D(A)\subset X\to X$ which satisfy
  1. (a)
    $\sigma(A)\subset S_{\omega}.$
  2. (b)
    For any $\omega<\mu<\pi$, there exists a positive constant $C_{\mu}$ such that
    $\|(z-A)^{-1}\|\leqslant C_{\mu}|z|^{\gamma},\quad\text{for any}\quad z\notin S_{\mu}.$
For simplicity we denote $\Theta_{\omega}^{\gamma}$ instead of $\Theta_{\omega}^{\gamma}(X)$. A linear operator $A$ will be called an almost sectorial operator in $X$ if $A\in\Theta_{\omega}^{\gamma}.$ Note that $0\in\rho(A)$ for any $A\in\Theta_{\omega}^{\gamma}.$ Also, we have that $A$ is injective [14, Remark 2.2]. It is important to recall that operators in the class $\Theta_{\omega}^{\gamma}$ have the possibility of having non-dense domain and/or range. This latter feature gives a different panorama with respect to classical results where it is always assumed dense domains [13, 16].
Let us now recall some necessary definitions and results that will be used through the whole paper.
Below we always denote by $\Gamma_{\theta}$ $(0<\theta<\pi)$ the path
$\{\rho e^{-i\theta}:\rho>0\}\cup\{\rho e^{i\theta}:\rho>0\}$
oriented such that the sector $S_{\theta}^{0}$ lies to the left of $\Gamma_{\theta}.$
First, we take $0<\alpha<\frac{\pi}{2\omega}$ and $t\in S_{\frac{\pi}{2}-\alpha\omega}^{0}$, with $\omega<\mu<\text{min}\left\{\pi,\frac{\pi-2|\text{arg}\,t|}{2\alpha}\right\}.$ Suppose that the function $g_{\alpha,t}(z)=e^{-tz^{\alpha}}$ is defined on $z\in S_{\mu}^{0}.$ Assume also that $A\in\Theta_{\omega}^{\gamma}.$ So, by [14, Lemma 2.13], we have that
$\mathscr{T}_{\alpha}(t)=\frac{1}{2\pi}\int_{\Gamma_{\theta}}e^{-tz^{\alpha}}(z-A)^{-1}{\rm d}z,\quad\omega<\theta<\mu,$
is a bounded linear operator in $X.$ In particular, the above holds for $0<\alpha\leqslant 1$ Note that the above expression for $\mathscr{T}_{\alpha}(t)$ could be also denoted by $e^{-tA^{\alpha}}.$ The latter notation is frequently used for strongly continuous semigroups generated by $A^{\alpha}$.
Next the definition about the analytic semigroups of growth order.
Definition 2.2.
Let $20<\mu<\pi/$ 2 and $\kappa>0.$ A family $\{\mathscr{T}(t):t\in S_{\mu}^{0}\}$ is said to be an analytic semigroup of growth order $\kappa$ if the following conditions hold:
  1. (1)
    $\mathscr{T}(t+s)=\mathscr{T}(t)\mathscr{T}(s)$ for any $t,s\in S_{\mu}^{0}.$
  2. (2)
    The mapping $t\to\mathscr{T}(t)$ is analytic in $S_{\mu}^{0}.$
  3. (3)
    There exists a positive constant $C$ such that
    $\|\mathscr{T}(t)\|\leqslant Ct^{-\kappa},\quad\text{for any}\quad t>0.$
  4. (4)
    If $\mathscr{T}(t)x=0$ for some $t\in S_{\mu}^{0}$, then $x=0.$
The above definition has two difference with respect to the one given by Da Prato. First, the set $X_{0}=\bigcup_{t>0}\mathscr{T}(t)X$ not need to be dense in $X.$ Second, the strong continuity of the mapping $t\to\mathscr{T}(t)$ for $t>0$ is replaced by condition (2).
Now we remember some properties of the operators families $\{\mathscr{T}_{\alpha}(t):t\in S_{\frac{\pi}{2}-\alpha\omega}\}$ stated in [14, Theorem 3.9] that will be used implicitly in some of our proofs.
Theorem 2.3.
Suppose that $A\in\Theta_{\omega}^{\gamma}$ for some $0<\omega<\frac{\pi}{2}$ and $0<\alpha<\frac{\pi}{2\omega}.$ Then the family $\{\mathscr{T}_{\alpha}(t):t\in S_{\frac{\pi}{2}-\alpha\omega}^{0}\}$ is an analytic semigroup of growth order $\frac{1+\gamma}{\alpha}.$ So, the following assertions are true.
  1. (1)
    $\mathscr{T}_{\alpha}(t+s)=\mathscr{T}_{\alpha}(t)\mathscr{T}_{\alpha}(s)$ for any $t,s\in S_{\frac{\pi}{2}-\alpha\omega}.$
  2. (2)
    There exists a positive constant $C(\gamma,\alpha)$ such that
    $\|\mathscr{T}_{\alpha}(t)\|\leqslant Ct^{-\frac{\gamma+1}{\alpha}},\quad\text{for any}\quad t>0.$
  3. (3)
    The range $R(\mathscr{T}_{\alpha}(t))$ of $\mathscr{T}_{\alpha}(t)$ with $t\in S_{\frac{\pi}{2}-\alpha\omega}^{0}$, is contained in $D(A^{\infty})$. In particular, $R(\mathscr{T}_{\alpha}(t))\subset D(A^{\beta})$ for all $\beta\in\mathbb{C}$ with $\text{Re}\,\beta>0,$ and
    $A^{\beta}\mathscr{T}_{\alpha}(t)x=\frac{1}{2\pi i}\int_{\Gamma_{\theta}}z^{\beta}e^{-tz^{\alpha}}(z-A)^{-1}x\,{\rm d}z,\quad\text{for all}\quad x\in X.$
    Also, there is a positive constant $C(\gamma,\alpha,\beta)$ such that
    $\|A^{\beta}\mathscr{T}_{\alpha}(t)\|\leqslant t^{-\frac{\gamma+1+{\rm Re}\,\beta}{\alpha}},\quad\text{for all}\quad t>0.$
  4. (4)
    The function $t\to\mathscr{T}_{\alpha}(t)$ is analytic in $S_{\frac{\pi}{2}-\alpha\omega}^{0}$ and
    $\frac{{\rm d}^{k}}{{\rm d}t^{k}}\mathscr{T}_{\alpha}(t)=(-1)^{k}A^{k\alpha}\mathscr{T}_{\alpha}(t),\quad\text{for all}\quad t\in S_{\frac{\pi}{2}-\alpha\omega}^{0}.$
  5. (5)
    Let $\Omega_{\alpha}=\left\{x\in X:\displaystyle\lim_{t\to 0,\,t>0}\mathscr{T}_{\alpha}(t)x=x\right\}$ be the continuity set of $\mathscr{T}_{\alpha}(\cdot)$. If $\beta>1+\gamma$ then $D(A^{\beta})\subset\Omega_{\alpha}.$
Abstract evolution equations. Suppose that $A\in\Theta_{\omega}^{\gamma}$ with $20<\omega<\pi/$ 2 and $-1<\gamma<0.$ As usual with this type of operators, we do not request density of the domain of $A.$
Consider the following abstract Cauchy problem:
(2.1) $\begin{split}\partial_{t}u:=u_{t}(t)&=A^{\alpha}u(t),\quad 0<t<T,\quad 0<\alpha<\frac{\pi}{2\omega},\\ u(0)&=u_{0}\in X.\end{split}$
By a classical solution of problem (2.1), we mean a function $u\in C\big{(}(0,T);X\big{)}\cap C^{1}\big{(}(0,T);X\big{)}$ which takes values in $D(A^{\alpha})$ for any $0<t<T$ and satisfies (2.1). It is important to highlight that if $u_{0}\in D(A^{\alpha})$ then the classical solution is continuous also in $t=0$ whenever $\alpha>1+\gamma$ (Theorem 2.3, item (5)).
It can be proved (see e.g. [14, Chapter 4]) that equation (2.1) has a unique classical solution given by
(2.2) $u(t)=\mathscr{T}_{\alpha}(t)u_{0},\quad\text{for any}\quad 0<t<T.$
3. Direct abstract Cauchy problem
In this section, we discuss the direct abstract Cauchy problem with a time-dependent variable coefficient $\phi$ by using an almost sectorial operator in the class $\Theta_{\omega}^{\gamma}$. The time-dependent function $\phi:[0,+\infty)\to[0,+\infty)$ is continuous with $\phi(t)>0$ for $t>0$.
We consider the abstract Cauchy problem in a complex Banach space $X$ with the coefficient $\phi(t)$ and $A\in\Theta_{\omega}^{\gamma}$ $\big{(}0<\omega<\pi/2\big{)}$ as follows:
(3.1) $\begin{split}u_{t}(t)&=\phi(t)A^{\alpha}u(t),\quad 0<t<T,\quad 0<\alpha<\frac{\pi}{2\omega},\\ u(0)&=u_{0}\in X.\end{split}$
Below we prove that the solution operator $\mathscr{T}_{\alpha,\phi}(t)$ of problem (3.1) has an explicit representation in terms of the coefficient $\phi(t)$ and the analytic semigroup $\mathscr{T}_{\alpha}(t)$ of growth order $\frac{\gamma+1}{\alpha}$.
Theorem 3.1.
Let $X$ be a complex Banach space. Then the solution operator of problem (3.1) is given by:
(3.2) $\mathscr{T}_{\alpha,\phi}(t)=\mathscr{T}_{\alpha}\left\{\left(\int_{0}^{t}\phi(s){\rm d}s\right)\right\},\quad 0<t<T.$
Proof.
Since $\phi$ is a continuous function such that $\phi(t)>0$ for any $t>0$, we have that $\eta(t)=\int_{0}^{t}\phi(r){\rm d}r$ for $t\geqslant 0$ is a bijection. Hence, there exits a unique function $\zeta:[0,+\infty)\to[0,+\infty)$ such that
(3.3) $\begin{split}\zeta(\eta(t))&=t,\quad t\geqslant 0,\\ \eta(\zeta(s))&=s,\quad s\geqslant 0.\end{split}$
Moreover, we get that $\partial_{s}\zeta(s)=\frac{1}{\phi(\zeta(s))}$ for any $s\geqslant 0$. So, we compute
$\begin{split}\frac{\rm d}{{\rm d}s}\big{(}u(\zeta(s))\big{)}&=\partial_{s}u(\zeta(s))\partial_{s}\zeta(s)=\big{(}\phi(\zeta(s))A^{\alpha}u(\zeta(s))\big{)}\frac{1}{\phi(\zeta(s))}=A^{\alpha}u(\zeta(s)),\end{split}$
where $u(\zeta(0))=u(0)=u_{0},$ for any $u$ satisfying equation (3.1). By using the uniqueness of the solution (2.2) of problem (2.1), we obtain
(3.4) $(u\circ\zeta)(s)=\mathscr{T}_{\alpha}(s)u_{0}\quad\text{for any}\quad s\geqslant 0.$
With $s=\eta(t)$, expression in (3.4) can be rewritten as
$u(t)=u(\zeta(\eta(t)))=\mathscr{T}_{\alpha}\big{(}\eta(t)\big{)}u_{0}=\mathscr{T}_{\alpha}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}u_{0},$
where $u(t)$ is the solution operator of (3.1). ∎
Let us now mention, in particular, the classical and important case $\alpha=1.$ Thus, suppose that:
(3.5) $\begin{split}u_{t}(t)&=\phi(t)Au(t),\quad 0<t<T,\\ u(0)&=u_{0}\in X,\end{split}$
where $A\in\Theta_{\omega}^{\gamma}$ with $20<\omega<\pi/$ 2 . Below we denote by $\mathscr{T}(t),$ $\mathscr{T}_{\phi}(t)$, the analytic semigroup $\mathscr{T}_{1}(t)$, and $\mathscr{T}_{1,\phi}(t)$ respectively.
Corollary 3.2.
Let $X$ be a complex Banach space. Then the solution operator of problem (3.5) can be obtained as:
(3.6) $\mathscr{T}_{\phi}(t)=\mathscr{T}\left\{\left(\int_{0}^{t}\phi(s){\rm d}s\right)\right\},\quad 0<t<T.$
4. Inverse abstract Cauchy problems
We now use the representation of the solution operator of the (direct) abstract Cauchy problem (found in the previous Section) to find the time-dependent coefficient for the inverse abstract Cauchy problem. The method is based in the consideration of two (direct) abstract Cauchy problems with the same time-dependent coefficient.
We start by establishing a useful property that involves the analytic semigroup of growth order and a fractional power of an operator. Let $A\in\Theta_{\omega}^{\gamma}$ for some $20<\omega<\pi/$ 2 . From [14, Theorem 2.5] and Theorem 2.3 (item (3)), we get that
(4.1) $A^{\alpha}\mathscr{T}_{\alpha}(t)x=\mathscr{T}_{\alpha}(t)A^{\alpha}x,\quad t\in S^{0}_{\frac{\pi}{2}-\alpha\omega},\quad x\in D(A^{\alpha}),\quad 0<\alpha<\frac{\pi}{2\omega},$
since the function $z^{\alpha}e^{-tz^{\alpha}}$ is in the class $\mathcal{F}_{0}^{\gamma}(S_{\mu}^{0})$ $\big{(}\omega<\theta<\mu<\text{min}\left\{\pi,\frac{\pi}{2\alpha}\right\}\big{)}.$ More details about it can be found in [14, Section 2].
For $A\in\Theta_{\omega}^{\gamma}$ $(0<\omega<\pi/2)$ and $\phi:[0,+\infty)\to[0,+\infty)$ being a continuous function with $\phi(t)>0$ for $t>0$, we consider the following two abstract Cauchy problems with the variable coefficient $\phi$:
(4.2) $\begin{split}u_{t}(t)&=\phi(t)A^{\alpha}u(t),\quad 0<t<T,\quad 0<\alpha<\frac{\pi}{2\omega},\\ u(0)&=u_{0}\in D(A^{\alpha}),\end{split}$
and
(4.3) $\begin{split}v_{t}(t)&=\phi(t)A^{\alpha}v(t),\quad 0<t<T,\quad 0<\alpha<\frac{\pi}{2\omega},\\ v(0)&=A^{\alpha}u_{0}\in D(A^{\alpha}).\end{split}$
By Theorem 3.1, the solution operators of problems (4.2) and (4.3) are given respectively by
(4.4) $u(t)=\mathscr{T}_{\alpha}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}u_{0},\quad\text{and}\quad v(t)=\mathscr{T}_{\alpha}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}A^{\alpha}u_{0}.$
Let us now fix an observation point $q$ for the following two time dependent quantities:
$h_{1}(t):=u(t,q),\quad h_{2}(t):=v(t,q),\quad 0<t<T.$
Next, we establish the main result of this section as follows.
Theorem 4.1.
Let $X$ be a complex Banach space. Then there exists a unique function $\phi(t)$ of the inverse abstract Cauchy problem (4.2) (and (4.3)) given by
$\phi(t)=\frac{\partial_{t}h_{1}(t)}{h_{2}(t)},\quad\text{for any}\quad 0<t<T.$
Proof.
By using (4.4), Theorem 2.3 (item (4)) and property (4.1) we obtain
$\partial_{t}h_{1}(t)=\partial_{t}u(t,q)$ $=\partial_{t}\mathscr{T}_{\alpha}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}u_{0}(t,q)$
$=\phi(t)A^{\alpha}\mathscr{T}_{\alpha}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}u_{0}(t,q)$
$=\phi(t)\mathscr{T}_{\alpha}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}A^{\alpha}u_{0}(t,q)$
$=\phi(t)v(t,q)=\phi(t)h_{2}(t),$
which completes the proof. ∎
Again, it is worthy to highlight the particular case $\alpha=1.$ Hence, let us study the following two abstract Cauchy problems with the variable coefficient $\phi$:
(4.5) $\begin{split}u_{t}(t)&=\phi(t)Au(t),\quad 0<t<T,\\ u(0)&=u_{0}\in D(A),\end{split}$
and
(4.6) $\begin{split}v_{t}(t)&=\phi(t)Av(t),\quad 0<t<T,\\ v(0)&=Au_{0}\in D(A),\end{split}$
where $A\in\Theta_{\omega}^{\gamma}$ $(0<\omega<\pi/2)$ and $\phi:[0,+\infty)\to[0,+\infty)$ is a continuous function with $\phi(t)>0$ for $t>0.$ The solution operators of problems (4.5) and (4.6) are given respectively by
$u(t)=\mathscr{T}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}u_{0},\quad\text{and}\quad v(t)=\mathscr{T}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}Au_{0}.$
Here, for the fixed point $q$, we denote by
$\widetilde{h_{1}}(t)=\mathscr{T}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}u_{0}(t,q)\quad\text{and}\quad\widetilde{h_{1}}(t)=\mathscr{T}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}Au_{0}(t,q).$
Corollary 4.2.
Let $X$ be a complex Banach space. Assume $\phi:[0,+\infty)\to[0,+\infty)$ is a continuous function such that $\phi(t)>0$ for $t>0$. Then there exists a unique function $\phi(t)$ of the inverse abstract Cauchy problem (4.5) (and (4.6)) constructed as
$\phi(t)=\frac{\partial_{t}\widetilde{h_{1}}(t)}{\widetilde{h_{2}}(t)},\quad\text{for any}\quad 0<t<T.$
5. Non-autonomous Cauchy problems
Suppose now that our linear operator $A$ depends also of a time-variable $\tau\geqslant 0$. In this case, the operator is usually denote by $A(\tau).$ By doing an analog of the classical definition, in [3, 8], the family of all linear closed operators $A(\tau):D(A(\tau))\subset X\to X$ which satisfy both conditions of Definition 2.1 with $A(\tau)$ instead of $A$, it is called an almost sectorial operator and it is again denoted by $\Theta_{\omega}^{\gamma}(X).$
Having in mind the study made in [1, 14], in [3, 8], it is then constructed and sometimes proved similarly many basic results for the case of $A(\tau).$ In this regard, we have that the semigroup $\{T_{A(\tau)}(t)\}_{t\in S_{\frac{\pi}{2}-\omega}}^{0}$ $(\omega<\theta<\mu<\frac{\pi}{2}-|\text{arg}\,t|,\,\text{for each}\,\tau\geqslant 0)$ of growth order $1+\gamma$ is given by
$T_{A(\tau)}(t)=\frac{1}{2\pi i}\int_{\Gamma_{\theta}}e^{-tz}(z-A(\tau))^{-1}x\,{\rm d}z,\quad t\in S_{\frac{\pi}{2}-\omega}^{0},$
and it is a bounded linear operator in $X.$ Here, the semigroup of growth order $1+\gamma$ means: i) $T_{A(\tau)}(0)=I;$ ii) $T_{A(\tau)}(t+s)=T_{A(\tau)}(t)T_{A(\tau)}(s)$ for each $t,s>0;$ iii) $T_{A(\tau)}(t)x=0$ implies $x=0$ for $t>0;$ iv) $\lim_{t\to 0}t^{1+\gamma}T_{A(\tau)}(t)$ is bounded.
Consider the following abstract Cauchy problem:
(5.1) $\begin{split}u_{t}(t)&=A(\tau)u(t),\quad 0<t<T,\quad\tau\geqslant 0,\\ u(0)&=u_{0}\in X.\end{split}$
By a classical solution of problem (5.1), we mean a function $u\in C\big{(}(0,T);X\big{)}\cap C^{1}\big{(}(0,T);X\big{)}$ that takes values in $D(A(\tau))$ for any $0<t<T$ and satisfies (5.1).
It is known that (see e.g. [3, Lemma 2.3]) that equation (5.1) has a unique classical solution given by
(5.2) $u(t)=T_{A(\tau)}(t)u_{0},\quad\text{for any}\quad 0<t<T,\quad\tau\geqslant 0.$
5.1. Direct abstract Cauchy problem
We study the abstract Cauchy problem with the coefficient $\phi(t)$ and $A(\tau)\in\Theta_{\omega}^{\gamma}$ $\big{(}0<\omega<\pi/2,\,\tau\geqslant 0\big{)}$ as follows:
(5.3) $\begin{split}u_{t}(t)&=\phi(t)A(\tau)u(t),\quad 0<t<T,\quad\tau\geqslant 0,\\ u(0)&=u_{0}\in X.\end{split}$
The proof of the next statement follows the same calculations and steps of the proof of Theorem 3.1. Therefore, we omit its proof and leave it to an interested reader.
Theorem 5.1.
Let $X$ be a complex Banach space. Then the solution operator of problem (5.3) is expressed as:
(5.4) $T_{A(\tau),\phi}(t)=T_{A(\tau)}\left\{\left(\int_{0}^{t}\phi(s){\rm d}s\right)\right\},\quad 0<t<T,\quad\tau\geqslant 0.$
5.2. Inverse abstract Cauchy problems
We start by establishing a property that involves the semigroup $T_{A(\tau),\phi}(t)$ of growth order $1+\gamma$ and the operator $A(\tau)$. Let $A(\tau)\in\Theta_{\omega}^{\gamma}$ for some $20<\omega<\pi/$ 2 and $\tau\geqslant 0.$ From [3, Pages 22-24] (or [8, Lemma 2.7 (ii)]), we know that
(5.5) $A(\tau)T_{A(\tau),\phi}(t)x=T_{A(\tau),\phi}(t)A(\tau)x,\quad t\in S^{0}_{\frac{\pi}{2}-\omega},\quad x\in D(A(\tau)),\quad\tau\geqslant 0.$
For $A(\tau)\in\Theta_{\omega}^{\gamma}$ $(0<\omega<\pi/2,\,\tau\geqslant 0)$ and $\phi:[0,+\infty)\to[0,+\infty)$ being a continuous function with $\phi(t)>0$ for $t>0$, we consider the following two abstract Cauchy problems with the variable coefficient $\phi$:
(5.6) $\begin{split}u_{t}(t)&=\phi(t)A(\tau)u(t),\quad 0<t<T,\quad\tau\geqslant 0,\\ u(0)&=u_{0}\in D(A(\tau)),\end{split}$
and
(5.7) $\begin{split}v_{t}(t)&=\phi(t)A(\tau)v(t),\quad 0<t<T,\quad\tau\geqslant 0,\\ v(0)&=A(\tau)u_{0}\in D(A(\tau)).\end{split}$
By Theorem 5.1, the solution operators of problems (5.6) and (5.7) are represented by
(5.8) $u(t)=T_{A(\tau)}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}u_{0},\,\,\text{and}\,\,v(t)=T_{A(\tau)}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}A(\tau)u_{0}.$
Let us now fix an observation point $q$ for the following two time dependent quantities:
$k_{1}(t):=u(t,q),\quad k_{2}(t):=v(t,q),\quad 0<t<T.$
Next, we establish the main result of this subsection.
Theorem 5.2.
Let $X$ be a complex Banach space. Then there exists a unique function $\phi(t)$ of the inverse abstract Cauchy problem (5.6) (and (5.7)) given by
$\phi(t)=\frac{\partial_{t}k_{1}(t)}{k_{2}(t)},\quad\text{for any}\quad 0<t<T.$
Proof.
By using (5.8), [3, Lemma 2.3] (or [8, Lemma 2.7 (i)]) and property (5.5) it follows that
$\partial_{t}k_{1}(t)=\partial_{t}u(t,q)$ $=\partial_{t}T_{A(\tau)}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}u_{0}(t,q)$
$=\phi(t)A(\tau)T_{A(\tau)}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}u_{0}(t,q)$
$=\phi(t)T_{A(\tau)}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}A(\tau)u_{0}(t,q)$
$=\phi(t)v(t,q)=\phi(t)k_{2}(t),$
that finishes the proof. ∎
Remark 5.3.
Note that equations in this section depend on operators $A(\tau)$ for some $\tau\geqslant 0.$ Therefore, we study, in principle, an evolution equation with respect to the time-variable $t$ and an almost sectorial operator $A(\tau).$ The case $A(t)$ (we mean e.g. $u_{t}(t)=A(t)u(t);$ $u(\tau)=u_{0};$ $\tau>t$) is in a sense different since the solution operator is associated to a new notion of linear evolution process introduced in [3]. Also, the case of fractional powers of $A(\tau)$ is not well studied yet. Therefore, these two cases will be analyzed somewhere else.
Conflict of interest
The authors declare that there are no conflicts of interest.